Show that velocity gradient = frequency.
1
Expert's answer
2019-04-02T10:45:13-0400

We can show that dimension of the velocity gradient is equal to dimension of frequency.


"\dim \frac{dv}{dx}=\frac{\dim v}{\dim x}=\frac{LT^{-1}}{L}=T^{-1}"

"\dim f =T^{-1}"

Thus, the dimension of the velocity gradient is equal to dimension of frequency.